A 17 kg block is pulled on a horizontal surface by a force of 205 N that is inclined 29° above the horizontal. If the coefficient of friction is 0.32 and the distance covered by the block was 6.25 meters. What is the work done by the friction force?


1
Expert's answer
2021-01-25T15:59:11-0500
"W_f=F_fs\\W_f=\mu Ns=\mu (mg-F\sin{29})s\\W_f=0.32 ((17)(9.8)-205\sin{29})(6.25)\\W_f=134\ J"